OpenStudy (anonymous):

integral dx/x^2-x+2 by trigonometric substitution... completing the square i got (x-(1/2))^2+7/4

OpenStudy (mathmale):

A bit messy, isn't this, Dan? I agree with your completion of the square. Now note that you could use the substitution u=x-1/2, ending up with \[\int\limits_{}^{}\frac{ du }{ u ^{2}+(\frac{ \sqrt{7} }{ 2 })^{2} }\]

OpenStudy (mathmale):

You with me so far?

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

Then, since we're asked to use trig subst., let u=\[u=\frac{ \sqrt{7} }{ 2 }\tan \theta\]

OpenStudy (mathmale):

\[u ^{2}=\frac{ 7 }{ 4 }\tan ^{2}\theta;du=\frac{ \sqrt{7} }{ 4 }\sec ^{2}\theta d \theta\]

OpenStudy (mathmale):

What would you do next? Have u used this particular trig subst. before?

OpenStudy (mathmale):

\[u ^{2}=\frac{ 7 }{ 4 }\tan ^{2}\theta;du=\frac{ \sqrt{7} }{ 2 }\sec ^{2}\theta d \theta\]

OpenStudy (anonymous):

yes, i used it, thanks a lot i appreciate it.

OpenStudy (mathmale):

My great pleasure. Obviously, once you've integrated with respect to theta, you'll have to work backwards and convert your answer in terms of theta to terms of (x-1/2) and then in terms of x. OK with that?

OpenStudy (mathmale):

Just let me know if further questions arise. All the best to you!

OpenStudy (anonymous):

yep, ty